OpenStudy (anonymous):

If f(x) = 3x + 2, find f(2).

OpenStudy (anonymous):

wherever you see an 'x', substitute with 2 and evaluate

OpenStudy (anonymous):

2=3x+2 0=3x 0=x

OpenStudy (anonymous):

@lawrenceh123 nope that's not the way; @AlexPR787 : f(2) = (3*2)+2 f(2) = 6+2 = 8

OpenStudy (anonymous):

Oh yeah, i was trying to find x, lol